The sum of three integers is 27. The second is 3 more than the first, and the third is two more than twice the second integer. Find the integers
Answer by jorel555(1290)

Let m,n,and p be your integers. Then
m+n+p=27
n=m+3
p=2n+2; or 2m+8
So:
m+m+3+2m+8=27
4m=16
m=4
n=7
p=16!!!!!!!!!!!!!!!